The time in hours that Davin plays on his games machine each day is normally distributed with mean 3.5 and standard deviation 0.9.
- (a)Find the probability that on a randomly chosen day Davin plays on his games machine for more than 4.2 hours.[3]
- (b)On 90% of days Davin plays on his games machine for more than t hours. Find the value of t.[3]
- (c)Calculate an estimate for the number of days in a year (365 days) on which Davin plays on his games machine for between 2.8 and 4.2 hours.[3]
Mark scheme
Marking points from the official mark scheme (Oct/Nov 2020). The mark or mark type is shown in brackets.
- [M1] Using ± standardisation formula, no √σ or σ², continuity correction
- [M1] Appropriate area Φ, from standardisation formula P(z>...) in final solution
- [A1] 0.218
- [B1] ±1.282 seen (critical value)
- [M1] An equation using standardisation formula with a z-value, condone √σ, σ² and continuity correction
- [A1] t = 2.35 (AWRT, only dependent on M mark)
- [B1FT] P(2.8 < X < 4.2) = 1 – 2 × their 5(a) OR P(2.8 < X < 4.2) = 2(Φ(0.7778) - 0.5) OR P(2.8 < X < 4.2) = Φ(0.7778) - (1 - Φ(0.7778)) = 0.5636. Accept unevaluated probability. OE. Accept 0.564.
- [M1] 365 × their p
- [A1FT] Accept 205 or 206, not 205.0 or 206.0, no approximation/rounding stated. FT must be an integer value.
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →